OpenStudy (anonymous):

Which of the following is a solution to sqrtx+10+3=8

OpenStudy (anonymous):

sqrtx = 8 - 3 - 10 sqrtx = -5 x = 25

OpenStudy (anonymous):

@i0iz0gangster

OpenStudy (anonymous):

you miswrote the question..

OpenStudy (anonymous):

\[\sqrt{x+10}+3=8\]

OpenStudy (anonymous):

a.x=15 b.x=5 c.x=-5 d.x=-15

OpenStudy (anonymous):

@i0iz0gangster

OpenStudy (i0iz0gangster):

I keep getting 6 man I don't think am be much help plus am taking a test too so am be abit busy

OpenStudy (anonymous):

oh sorry thank for looking at it for me

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

Ok, simple enough!

OpenStudy (anonymous):

really?:\

OpenStudy (austinl):

\[\sqrt{x+10}+3=8\] You would subtract the 3 over first to get \[\sqrt{x+10}=5\] Then you would square both sides. \[(\sqrt{x+10})^2=(5)^2\] \[x+10=25\] Subtract the 10 over \[x=15\]

OpenStudy (anonymous):

thank you so much i am going to write that down for my test tomorrow.

OpenStudy (austinl):

Haha, no problem!